# LLM-Based Detection of Tangled Code Changes for Higher-Quality Method-Level Bug Datasets

### v2

@@ Abstract.
Tangled code changes, commits that conflate unrelated modifications such as bug fixes, refactorings, and enhancements, introduce significant noise into bug datasets and adversely affect the performance of bug prediction models. Addressing this issue at a fine-grained, method-level granularity remains underexplored. This is critical to address, as recent bug prediction models, driven by practitioner demand, are increasingly focusing on finer granularity rather than traditional class- or file-level predictions. This study investigates the utility of Large Language Models (LLMs) for detecting tangled code changes by leveraging both commit messages and method-level code diffs. We formulate the problem as a binary classification task and evaluate multiple prompting strategies, including zero-shot, few-shot, and chain-of-thought prompting, using state-of-the-art proprietary LLMs such as GPT-4o and Gemini-2.0-Flash.
Our results demonstrate that combining commit messages with code diffs significantly enhances model performance, with the combined few-shot and chain-of-thought prompting achieving an F1-score of 0.88. Additionally, we explore machine learning models trained on LLM-generated embeddings, where a multi-layer perceptron classifier achieves superior performance (F1-score: 0.906, MCC: 0.807). Applying our approach to 49 open-source projects improves the distributional separability of code metrics between buggy and non-buggy methods, demonstrating the promise of LLMs for method-level commit untangling and potentially contributing to improving the accuracy of future bug prediction models.
@@ 1. Introduction
Software maintenance often exceeds the cost of initial development (Börstler and Paech, 2016), with bug detection and resolution alone consuming 50–70% of total development expenses (Wang et al., 2023). Undetected bugs can result in cascading failures, reduced developer productivity, and increased technical debt (christine fisher, 2020; sixsentix, 2024). Consequently, the research community has extensively investigated bug prediction techniques to facilitate timely and targeted maintenance interventions (Chowdhury et al., 2024; Pascarella et al., 2020; Wattanakriengkrai et al., 2022; Giger et al., 2012; Rahman and Roy, 2017). Early bug prediction models focused on class- or file-level granularity (Zimmermann et al., 2007; Alsolai et al., 2018; Basili et al., 1996; Bell et al., 2011; Gil and Lalouche, 2017; Zhou et al., 2010; Sun et al., 2025; Yin et al., 2025). However, these models are less helpful in practice, as only a small fraction of code in large units is typically defective (Mo et al., 2022; Pascarella et al., 2020). Line-level models offer finer granularity but are prone to false positives due to coincidental line similarity (Chowdhury et al., 2024; Servant and Jones, 2017; Grund et al., 2021). This has driven growing interest in method-level bug prediction (Chowdhury et al., 2024; Mo et al., 2022; Ferenc et al., 2020; Pascarella et al., 2020; Shippey et al., 2016; Giger et al., 2012; Hata et al., 2012; Menzies et al., 2007), which aligns more closely with real-world debugging practices (Grund et al., 2021; Chowdhury et al., 2024). However, method-level bug prediction faces challenges including evaluation bias and noisy data (Pascarella et al., 2020; Chowdhury et al., 2024), and progress is constrained by the scarcity of clean, labeled datasets (Chowdhury et al., 2024; Pascarella et al., 2020).
[…]
@@
Given that, despite the demand, method-level bug prediction remains an open research challenge (Pascarella et al., 2020; Chowdhury et al., 2024), particularly due to the confounding effects of tangled changes. This study focuses on detecting tangled changes at the method-level by framing the task as a binary classification problem: determining whether a code change inside a method is related to a bug fix (Buggy) or not (NotBuggy) based on commit messages and code diffs. To tackle this, we explore the use of Large Language Models (LLMs), which have demonstrated strong capabilities in jointly analyzing natural language and source code (Abburi et al., 2023; Nam et al., 2024). LLMs provide advantages such as promptability, flexibility in input handling, and the ability to leverage contextual signals from both code and commit metadata. We evaluate several prompting strategies, including zero-shot, few-shot, and chain-of-thought prompting, using proprietary LLMs such as GPT-4o and Gemini-2.0-Flash. Additionally, we develop machine learning models leveraging embeddings generated by an LLM. Finally, we evaluate the impact of the LLM-based technique by measuring improvements in the distributional separability of code metrics between Buggy and NotBuggy methods through statistical analysis. In general, the contribution of the paper is centered around four research questions.
@@
We designed a classification-based approach that uses LLMs to detect tangled changes from method-level code diffs, using zero-shot prompting, with or without the corresponding commit messages. Through experiments on our curated gold dataset, we show that including the commit message significantly improves the performance across multiple LLM variants, achieving the highest F1-score of 0.877 using gpt-4o.
@@
To answer this question, we investigate the impact of advanced prompting strategies, such as few-shot learning, chain-of-thought reasoning, and a hybrid combination, on tangled change detection. Our analysis demonstrates that LLMs can reason more accurately about change semantics when guided with structured prompts and exemplar-based context. Among these, chain-of-thought with few-shot prompting yields the highest performance of 0.883 F1-score with the ability to detect complex and ambiguous input.
@@
We utilized LLM to generate embeddings from commit messages and code diffs, which served as input for embedding-based classifiers. These models exhibit strong predictive capabilities, outperforming the results obtained in RQ1 and RQ2. Specifically, a Multi-layer Perceptron classifier achieved the highest F1-score of 0.906, demonstrating the effectiveness of embedding-based representations for detecting tangled changes.
@@
To enable replication and extension, we share our data and code publicly. [^1] https://github.com/SQMLab/Tangled
@@ 2.1. Tangled Changes
Herzig and Zeller (Herzig and Zeller, 2013) first used the term Tangled Changes and reported that up to 15% of bug-fixing commits across several open-source Java projects contained tangled changes. Subsequently, Tao and Kim (Tao and Kim, 2015) found that tangled commits accounted for up to 29% of revisions in open-source projects. Herbold et al. (Herbold et al., 2022) further validated the high occurrence and detrimental effects, showing that only 22% to 38% of the lines in a bug fix commit contribute to bug fixes. Kochhar et al. (Kochhar et al., 2014) analyzed 100 bug-fixing commits to find out the implication of tangling on bug localization, showing that 28% of all files in bug fix commits do not have bugs.
[…]
Graph-based modeling has emerged as a powerful approach for capturing complex code relationships. Shen et al. (Shen et al., 2021) proposed SmartCommit, representing diff hunks as graph nodes enriched with semantic links (hard, soft, refactoring, cosmetic), and achieved a decomposition accuracy between 71-84%. Chen et al. (Chen et al., 2022) and Xu et al. (Xu et al., 2025) enhanced this approach by incorporating node attributes such as token content and control/data flow, achieving 7.8% and 8.2% higher accuracy than Flexeme, respectively.
@@
Although these methods have shown success at the file- or statement-level, the method-level remains largely unexplored, despite being preferred by practitioners and researchers (Pascarella et al., 2020; Chowdhury et al., 2024). Also, developing accurate bug prediction models at the method-level requires a clean and noise-free bug dataset, which is currently unavailable. Motivated by this gap, we investigate whether tangled changes can be detected from method-level diffs and their corresponding commit messages.
@@ 2.2. LLMs for Code Analysis & Classification
LLMs have shown strong performance in natural language understanding and source code analysis (Brown et al., 2020; Chen et al., 2021). Trained on large corpora of text and code, they generalize well to unseen inputs and often perform effectively in zero-shot settings (Kojima et al., [n. d.]). This versatility has fueled growing interest in using LLMs for text classification and code analysis (Ahmad et al., 2021; Feng et al., 2020). Building on these capabilities, our study explores LLMs for detecting tangled code changes by jointly analyzing commit messages and code diffs.
Prompt design is critical to optimizing LLM performance. Recent advances in prompt engineering have introduced zero-shot, few-shot, and chain-of-thought prompting (Wei et al., 2022; Brown et al., 2020; Wei et al., 2023). Zero-shot learning uses only task instructions, while few-shot prompting includes exemplar input-output pairs to guide responses, yielding improved performance across tasks (Brown et al., 2020). Chain-of-thought prompting enhances reasoning by encouraging step-wise problem decomposition (Wei et al., 2023). Motivated by their empirical success, we systematically evaluate these techniques for detecting tangled code changes.
LLMs are built on Transformer architectures, consisting of encoder and decoder modules (Vaswani et al., 2017). Encoders convert input text into high-dimensional embeddings that capture semantic relationships, making them effective for downstream classification (Petukhova et al., 2024; Keraghel et al., 2024). We extend this utility to software engineering by generating embeddings from commit messages and code diffs, and training machine learning models to classify them as tangled or untangled.
In summary, this study is motivated by three factors: (1) the lack of fine-grained research on the tangled changes at the method level, especially for bug prediction (Herzig et al., 2016; Chowdhury et al., 2024); (2) the proven capabilities of LLMs in analyzing textual and code artifacts via advanced prompting (Brown et al., 2020; Chen et al., 2021; Wei et al., 2022; Brown et al., 2020; Wei et al., 2023); and (3) the semantic richness of LLM-generated embeddings for classification (Petukhova et al., 2024; Keraghel et al., 2024). To the best of our knowledge, this is the first study to evaluate LLMs for detecting tangled code changes within bug-fix commits at the method level.
@@ 3. Methodology
Figure 1. The methodology for RQ1 and RQ2 consists of three parts: (a) dataset, (b) prompt preparation, and (c) LLM Query. The prompt setup for few-shot + chain-of-thought prompting consists of four instruction parts: (i) persona, (ii) task description, (iii) behavioral guidance, and (iv) output formatting. For Diff-only detection of RQ1, references to the commit message are excluded from everywhere. Other experiments use the full instructions, with the output formatting modified to require a single-word answer in non-chain-of-thought setups. Few-shot examples are included only in few-shot experiments. The dataset component also relates to RQ3; the remaining methodology for RQ3 is discussed later.
[…]
Overview of methodology.
@@ 3.1. Dataset
However, the dataset contains noise. In some cases, methods involved in a bug fix commit may not contribute to the actual fix, representing instances of tangled changes. We use this dataset to construct a curated gold dataset by removing such noise. The construction process involves a two-step sampling strategy, combining automated filtering with manual validation, to ensure high label quality and accurate representation of tangled change scenarios.
@@
Automated Approach. The dataset provided by Chowdhury et al. (Chowdhury et al., 2024) includes the total number of methods modified in each commit. In their study, they recommended that if a method is the only one modified in a bug fix commit, it can be confidently considered a bug-related change. We classify such code diffs as Buggy without ambiguity. Using this criterion, we collected 730 Buggy diffs. They also suggested that a method can be considered NotBuggy if it has never appeared in a bug fix commit throughout its lifetime, provided that it has had sufficient time to evolve. They recommended a threshold of two years. Based on this guideline and leveraging the large pool of eligible NotBuggy methods, we randomly selected 730 samples that are older than two years and have no history of being part of any bug fix commit. During this process, three duplicate diffs, likely caused by duplicated code across codebases, were identified and removed, resulting in 727 unique NotBuggy instances. Thus, the Automated step of our gold dataset construction produced a total of 1,457 method-level examples. We intentionally limited the number of NotBuggy samples, as querying proprietary LLMs incurs significant computational and financial costs. Including all such samples would have significantly increased the overall cost.
@@
Manual Approach. While the Automated approach ensured high confidence in labeling, it did not capture the complexity of real-world tangled change scenarios, where multiple unrelated modifications often occur within a single commit. To address this limitation, we introduced a second step focused on incorporating tangled examples. In this step, we extracted additional method-level diffs from bug-fix commits that modified multiple methods and conducted manual annotations to determine whether each method-level change was genuinely related to a bug fix (Buggy) or was incidental, such as refactorings or unrelated enhancements (NotBuggy). Each method was reviewed in conjunction with its associated commit message to evaluate semantic alignment with bug-fixing intent. The evaluation emphasized structural corrections, error handling logic, and domain-specific repairs. To promote annotation consistency and minimize errors, we developed a user interface that allowed side-by-side inspection of code diffs and their corresponding commit metadata. Using this Manual approach, we identified 166 tangled Buggy methods and 141 tangled NotBuggy methods.
Combining all steps, the gold dataset consists of 1,764 method-level change instances, each labeled as either Buggy or NotBuggy, based on their actual relevance to bug fixing activities. This dataset serves as a high-confidence benchmark for both the training and evaluation of our detection approaches, as well as a valuable resource for future research.

### v3

@@ Abstract.
Tangled code changes, commits that conflate unrelated modifications such as bug fixes, refactorings, and enhancements, introduce significant noise into bug datasets and adversely affect the performance of bug prediction models. Addressing this issue at a fine-grained, method-level granularity remains unexplored. This is critical to address, as recent bug prediction models, driven by practitioner demand, are increasingly focusing on finer granularity rather than traditional class- or file-level predictions. This study investigates the utility of Large Language Models (LLMs) for detecting tangled code changes by leveraging both commit messages and method-level code diffs. We formulate the problem as a binary classification task and evaluate multiple prompting strategies, including zero-shot, few-shot, and chain-of-thought prompting, using state-of-the-art proprietary LLMs such as GPT-5 and Gemini-2.0-Flash, and open-source models such as GPT-OSS-120B and CodeBERT.
Our results demonstrate that combining commit messages with code diffs significantly enhances model performance, with the combined few-shot and chain-of-thought prompting achieving an F1-score of 0.883. Additionally, we explore machine learning models trained on LLM-generated embeddings, where a multi-layer perceptron classifier achieves superior performance (F1-score: 0.906, MCC: 0.807). Applying our approach to 49 open-source projects improves the distributional separability of code metrics between buggy and non-buggy methods, demonstrating the promise of LLMs for method-level commit untangling and potentially contributing to improving the accuracy of future bug prediction models.
@@ 1. Introduction
Software maintenance often exceeds the cost of initial development (Börstler and Paech, 2016), with bug detection and resolution alone consuming 50–70% of total development expenses (wang_bugpre_2023). Undetected bugs can result in cascading failures, reduced developer productivity, and increased technical debt (christine fisher, 2020; noauthor_most_2024). Consequently, the research community has extensively investigated bug prediction techniques to facilitate timely and targeted maintenance interventions (Chowdhury et al., 2024; pascarella_performance_2020; wattanakriengkrai_predicting_2022; Giger et al., 2012; rahman_relationships_2017). Early bug prediction models focused on class- or file-level granularity (zimmermann_predicting_2007; Alsolai et al., 2018; Basili et al., 1996; Bell et al., 2011; Gil and Lalouche, 2017; sun_boosting_2025; yin_line-level_2025). However, these models are less helpful in practice, as only a small fraction of code in large units is typically defective (Mo et al., 2022; pascarella_performance_2020). Line-level models offer finer granularity but are prone to false positives due to coincidental line similarity (Chowdhury et al., 2024; servant_fuzzy_2017; Grund et al., 2021). This has driven growing interest in method-level bug prediction (Chowdhury et al., 2024; Mo et al., 2022; pascarella_performance_2020; Shippey; Giger et al., 2012; Hata et al., 2012; Menzies et al., 2007), which aligns more closely with real-world debugging practices (Grund et al., 2021; Chowdhury et al., 2024). However, method-level bug prediction faces challenges including evaluation bias and noisy data (pascarella_performance_2020; Chowdhury et al., 2024), and progress is constrained by the scarcity of clean, labeled datasets (Chowdhury et al., 2024; pascarella_performance_2020).
A key obstacle in building clean bug datasets is the presence of noise in training data (Bird et al., 2009), often caused by tangled code changes (Chowdhury et al., 2024; PârΩ, 2020; Herzig et al., 2016). While Version Control Systems like Git are intended for atomic commits, developers frequently combine unrelated tasks, such as bug fixes, refactorings, and enhancements, into a single commit (Herzig and Zeller, 2013; Dias et al., 2015; Kirinuki et al., 2014), violating best practices and introducing tangled changes. Researchers typically use commit messages to identify bug-fix commits, assuming all associated modifications are relevant. However, this assumption fails in the presence of tangled commits, leading to false positives in labels, which degrades dataset quality and impairs bug prediction performance (Herzig et al., 2016; Herbold et al., 2022).
To mitigate the impact of tangled changes, several prior works have proposed various untangling techniques ranging from heuristic-based program slicing (Muylaert and De Roover, 2018), lexical and data-flow based models (PârΩ, 2020), and graph-based learning using dependency and name flows (Li et al., 2022; xu_detecting_2025). While effective at the file or statement level, none of these methods have addressed the problem at the method-level granularity, leaving a gap that presents a unique opportunity for exploration. Method-level untangling cannot be achieved using file-level approaches, as a single file may contain multiple methods, only some of which are bug-related. Labeling all methods in a file as buggy is inappropriate and would produce numerous false positives. On the other hand, although statements can be aggregated to form method-level representations, determining whether an individual statement contributes to a bug fix requires understanding its broader context, including surrounding code, method semantics, and commit intent (tufano2019empirical; sintaha2023katana). For developers and models alike, the relevant context may span from within the method to the entire file or even across files (sintaha2023katana). For instance, in the refill function of FixedIntervalRateLimiter.java from Apache HBase Commit 84a5039  [^1] https://github.com/apache/hbase/commit/84a50393ee56d09abb68f54b44b64f5279bd33c9, the function signature modification appears to be a bug fix when viewed in isolation, as removing an argument typically alters the function’s logic; however, examining the full method reveals it is actually a refactoring change. Thus, simply applying a statement-level model repeatedly at the method-level would require redundant context analysis for each modified line, introducing unnecessary computational overhead. Therefore, distinct representations at different granularity are needed to effectively model bug-fix behavior while maintaining contextual relevance.
@@
Given that, this study focuses on detecting tangled changes at the method-level by framing the task as a binary classification problem: determining whether a method-level code change is related to a bug fix (Buggy) or not (NotBuggy) based on commit messages and code diffs. To tackle this, we explore the use of Large Language Models (LLMs), which have demonstrated strong capabilities in analyzing natural language and source code (Chew et al., 2023; Nam et al., 2024). In contrast to prior approaches in code analysis, classification, and untangling that rely on language-specific algorithms (Herzig et al., 2016; Guo and Song, 2017) or complex models requiring extensive training and feature engineering (zeng2024colare), LLM-based methods enable faster system development through their language-agnostic design. LLMs also provide advantages such as promptability, flexibility in input handling, and the ability to leverage contextual signals from both code and commit metadata.
We evaluate several prompting strategies using both proprietary and open-source LLMs. Additionally, we develop machine learning models using LLM generated embeddings. Finally, we evaluate the impact of the LLM-based technique by measuring improvements in the distributional separability of code metrics between Buggy and NotBuggy methods through statistical analysis. In general, the contribution of the paper is centered around four research questions.
@@
We designed a classification-based approach that uses LLMs to detect tangled changes from method-level code diffs, using zero-shot prompting, with or without the corresponding commit messages. Through experiments on our curated gold dataset, we show that including the commit message significantly improves the performance across multiple LLM variants, achieving the highest F1-score of 0.879 using gpt-5.
@@
We investigate the impact of advanced prompting strategies, such as few-shot, chain-of-thought, and a hybrid combination. Our analysis shows that LLMs can reason more accurately about change semantics when guided with structured prompts and examples. Among these, chain-of-thought + few-shot prompting with gpt-4o achieves the most balanced performance (F1-score: 0.883) and demonstrates superior capability in handling complex inputs.
@@
We utilized LLMs to generate embeddings from commit messages and code diffs, which served as input for embedding-based classifiers. One of the models exhibits strong predictive capabilities, outperforming the results obtained in RQ1 and RQ2. Specifically, a Multi-layer Perceptron classifier achieved the highest F1-score of 0.906, demonstrating the effectiveness of embedding-based representations for detecting tangled changes.
@@
To enable replication, we share our data and code publicly. [^2] https://github.com/SQMLab/Tangled
@@ 2.1. Tangled Changes
Herzig and Zeller (Herzig and Zeller, 2013) first used the term Tangled Changes and reported that up to 15% of bug-fixing commits across several open-source Java projects contained tangled changes. Subsequently, Tao and Kim (tao_partitioning_2015) later observed tangling in up to 29% of revisions. Herbold et al. (Herbold et al., 2022) confirmed its prevalence and impact, showing that only 22–38% of lines in bug-fix commits actually fix bugs. Kochhar et al. (Kochhar et al., 2014) further found that 28% of files in bug-fix commits contain no bugs.
The prevalence of tangled changes greatly affects the historical analysis of source code. By labeling all the modified methods in a bug-fixing commit as buggy methods, it harms the bug prediction models significantly (Herzig et al., 2016; Chowdhury et al., 2024). Consequently, researchers focused on detecting and untangling commits (Herzig and Zeller, 2013; tao_partitioning_2015; Guo and Song, 2017; Muylaert and De Roover, 2018; PârΩ, 2020).
Early techniques relied on static analysis and heuristics. Herzig et al. (Herzig and Zeller, 2013) introduced confidence voters based on file distance, change coupling, call graphs, and data dependencies to estimate task co-membership. Tao and Kim (tao_partitioning_2015) applied program slicing and pattern matching to cluster semantically related edits, achieving 69% agreement with manual decompositions. Guo and Song (Guo and Song, 2017) proposed CHGCUTTER, an interactive method using control and data dependencies to decompose composite changes while preserving syntactic correctness. Muylaert and De Roover (Muylaert and De Roover, 2018) applied program slicing on abstract syntax trees to untangle composite commits by grouping fine-grained code changes according to their dependence graph slices. Pârtachi et al. (PârΩ, 2020) introduced Flexeme, which overlays name flows on program dependency graphs using a $\delta$-NFG structure to better capture lexical cues. Using Agglomerative clustering for commit untangling they developed a tool Heddle, which achieved improved clustering accuracy (F1-score 0.81) and runtime efficiency.
Graph-based modeling has emerged as a powerful approach for capturing complex code relationships. Shen et al. (shen_smartcommit_2021) proposed SmartCommit, representing diff hunks as graph nodes enriched with semantic links (hard, soft, refactoring, cosmetic), and achieved a decomposition accuracy between 71-84%. Chen et al. (Chen et al., 2022) and Xu et al. (xu_detecting_2025) enhanced this approach by incorporating node attributes such as token content and control/data flow, achieving 7.8% and 8.2% higher accuracy than Flexeme, respectively.
@@
Although these methods have shown success at the file- or statement-level, the method-level remains unexplored, despite being preferred by practitioners and researchers in order to create a clean and noise-free method-level bug dataset (pascarella_performance_2020; Chowdhury et al., 2024). Motivated by this gap, we investigate whether tangled changes can be detected from method-level diffs and their corresponding commit messages.
@@ 2.2. LLMs for Code Analysis & Classification
LLMs have shown strong performance in natural language understanding and source code analysis (Brown et al., 2020; zheng2024few). Trained on large corpora of text and code, they generalize well to unseen inputs and often perform effectively in zero-shot settings (Kojima et al., 2022). This versatility has fueled growing interest in using LLMs for text classification and code analysis (Ahmad et al., 2021; Feng et al., 2020). Building on these capabilities, our study explores LLMs for detecting tangled code changes by jointly analyzing commit messages and code diffs.
Prompt design is critical to optimizing LLM performance. Recent advances in prompt engineering have introduced zero-shot, few-shot, and chain-of-thought prompting (wei_finetuned_2022; Brown et al., 2020; wei_chain--thought_2023). Zero-shot learning uses only task instructions, while few-shot prompting includes exemplar input-output pairs to guide responses, yielding improved performance across tasks (Brown et al., 2020). Chain-of-thought prompting enhances reasoning by encouraging step-wise problem decomposition (wei_chain--thought_2023). Motivated by their empirical success, we systematically evaluate these techniques for detecting tangled code changes.
LLMs are built on Transformer architectures, consisting of encoder and decoder modules (vaswani_attention_2017). Encoders convert input text into high-dimensional embeddings that capture semantic relationships, making them effective for downstream classification (petukhova_text_2024; Keraghel et al., 2024). We extend this utility to software engineering by generating embeddings from commit messages and code diffs, and training machine learning models to classify them as tangled or untangled.
LLM embeddings are also used in commit classification tasks (zeng2025first; zeng2024colare), which classifies commits into maintenance activity types. CodeBERT (Feng et al., 2020) has been used in COLARE (zeng2024colare) to classify commits into corrective, adaptive, and perfective categories by integrating hunk-level code representations with commit messages and file features. However, such models are build for commit-level classification which does not properly align with our method-level objectives.
In summary, this study is motivated by three factors: (1) the lack of fine-grained research on the tangled changes at the method level, especially for bug prediction (Herzig et al., 2016; Chowdhury et al., 2024); (2) the proven capabilities of LLMs in analyzing textual and code artifacts via advanced prompting (Brown et al., 2020; zheng2024few; wei_finetuned_2022; wei_chain--thought_2023); and (3) the semantic richness of LLM-generated embeddings for classification (petukhova_text_2024; Keraghel et al., 2024). To the best of our knowledge, this is the first study to evaluate LLMs for detecting tangled code changes within bug-fix commits at the method level.
@@ 3. Methodology
Figure 1. The methodology for RQ1 and RQ2 consists of three parts: (a) dataset, (b) prompt preparation, and (c) LLM Query. The prompt setup for few-shot + chain-of-thought prompting consists of four instruction parts: (i) persona, (ii) task description, (iii) behavioral guidance, and (iv) output formatting. For Diff-only detection of RQ1, references to the commit message are excluded from everywhere. Other experiments use the full instructions, with the output formatting modified to require a single-word answer in non-chain-of-thought setups. Few-shot examples are included only in few-shot experiments.:
@@ 3.1. Dataset
However, the dataset contains noise. In some cases, methods involved in a bug fix commit may not contribute to the actual fix, representing instances of tangled changes. We use this dataset to construct a curated gold dataset by removing such noise. The construction process involves a two-step sampling strategy, combining automated filtering and manual labeling, to ensure high quality and accurate representation of tangled change scenarios.
@@
Automated Approach. The dataset provided by Chowdhury et al. (Chowdhury et al., 2024) includes the total number of methods modified in each commit. They recommended that if a method is the only one modified in a bug fix commit, it can be confidently considered a bug-related change. We classify such code diffs as Buggy without ambiguity. Using this criterion, we collected 730 Buggy diffs. They also suggested that a method can be considered NotBuggy if it has never appeared in a bug fix commit throughout its lifetime. Based on this guideline and leveraging the large pool of eligible NotBuggy methods, we randomly selected 730 samples that have no history of being part of any bug fix commit. During this process, three duplicate diffs, likely caused by duplicated code across codebases, were identified and removed, resulting in 727 unique NotBuggy instances. Thus, the Automated step of our gold dataset construction produced a total of 1,457 method-level examples. We intentionally limited the number of NotBuggy samples, as querying proprietary LLMs incurs significant computational and financial costs. Including all such samples would have significantly increased the overall cost.
@@
Manual Approach. While the Automated approach ensured high confidence in labeling, it did not capture the complexity of real-world tangled change scenarios, where multiple unrelated modifications often occur within a single commit. To address this limitation, we introduced a second step focused on incorporating tangled examples. In this step, we extracted additional method-level diffs from bug-fix commits that modified multiple methods. The first author, a graduate student with over two years of industry experience in software engineering, manually labeled each method-level change to determine whether it was related to a bug fix (Buggy) or was incidental, such as refactorings or unrelated enhancements (NotBuggy) Each method was reviewed in conjunction with its associated commit message to evaluate semantic alignment with bug-fixing intent. The evaluation emphasized structural corrections, error handling logic, and domain-specific repairs. To promote annotation consistency and minimize errors, we developed a user interface that allowed side-by-side inspection of code diffs and their corresponding commit metadata. Using this manual approach, we identified 166 tangled Buggy methods and 141 tangled NotBuggy methods from 450 randomly selected commits in the dataset of Chowdhury et al. (Chowdhury et al., 2024).
Combining all steps, the gold dataset consists of 1,764 method-level change instances, each labeled as either Buggy or NotBuggy, based on their actual relevance to bug fixing activities. To evaluate labeling reliability, an independent validation was conducted by another graduate student in software engineering, who randomly picked 100 samples from the manually labeled portion of the dataset and labeled them independently. Some disagreements arose, mainly due to differing interpretations of what constitutes a bug. Specifically, the first labeler marked some quality-related fixes as NotBuggy, while the second considered them Buggy. Following a discussion with the last author, quality-fix-related changes were finalized as NotBuggy. The inter-rater agreement, measured using Cohen’s Kappa (Cohen, 1960), was 0.82, indicating strong consistency and confirming the reliability of the labeling process. Consequently, this dataset provides a robust benchmark for both training and evaluating our detection approaches, while also offering a valuable foundation for future research in this domain.